SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

DCH Run Log

[BaBar Run List] [Electronic Logbook]

Add entries to table any time conditions change that may have an effect on drift chamber data. Note that in some cases an entry appears for when the change occurred and separate entries for the first cosmic and/or colliding beam run after the change.

Last modified: 2 August 2000 by D. Coupal

Date Time Run Number Type Comment
29-JUL-99 18:04 Turned down SL6, outer quadrant to 1500v (effectively off).List of wires turned down.
29-JUL-99 18:10 7137 Cosmic?
29-JUL-99 20:31 7146 Colliding
06-AUG-99 14:00 7354 Colliding Water bubbler started
07-AUG-99 03:57 7379 Colliding Beginning of sharp drop in gain due to water level in gas
12-AUG-99 05:27 7725 Colliding Maximum gain drop (about 15 %)
14-AUG-99 09:19 7764 Colliding first colliding beams after drying out chamber
22-OCT-99 17:44 Set V1 to 1900v.  Turned SL5 bottom & outer, SL6 bottom and lowest 3 HV cards in SL6 outer down to V1=1200v.List of wires turned down.
22-OCT-99 20:53 9929 Cosmic
22-OCT-99 22:56 9931 Colliding
03-NOV-99       Re-segment the HV: see HN posting, List of wires turned down.
03-NOV-99 18:19 10224 Cosmic  
05-NOV-99 12:38 10256 Colliding  
27-JAN-00 17:48 11075-078 Cosmics Cosmic tests at lower voltage. See HN posting
04-APR-00 17:00 12626 Colliding First of two runs with sparsing set to TDC mode HN posting
04-APR-00 19:56 12628 Colliding Back to FADC sparsing mode.
07-APR-00 06:30 12662 Colliding First run to be processed with 45 micron stereo offset.
28-APR-00 17:03 13011 Colliding Single test run at 1960 volts. See HN posting
10-JUL-00 09:17 14462 Colliding First run at 1960 volts and TDC mode. See HN posting